OpenStudy (anonymous):

Randall and Augustine are financing $148,000 to purchase a condominium. They obtained a 15-year, fixed-rate loan with a rate of 4.95%. They have been given the option of purchasing up to five points to lower their rate to 4.73%. How much will the five points cost them?

OpenStudy (anonymous):

I got 7,326 but that isn't any of my answer choices

OpenStudy (anas.p):

Simple Interest formula is P*R*T wherein R is in decimals.( I mean 75% = 0.75) 148*0.0495*15=109.8 This is the interest they would have to pay 148*0.0473*15=105.0 109.8-105.0=4.8 Does that match anything.

OpenStudy (anonymous):

No , you used the wrong formula. $7,000.40 $325.60 $1,480.00 $7,400.00

OpenStudy (anas.p):

multiply it by thousand 4.8 * 1000 = 4800 sorry forgot to write that

OpenStudy (anas.p):

what formula did you use?

OpenStudy (anonymous):

I just followed an example and did 148000x0.05 = 7,400

OpenStudy (anonymous):

Huh I just want to know If i Got it right.